a family is buying capetibg for two rooms.One room measures 15 feet by 17.5 feet and the other is 17.5 feet by 20 feet.Which expression gives the total square eet of carpet that the family is buying?
F.35*35
G.32.5*37.5
H.17.5*35
I.15*37.5

Answer by checkley77(12844)

H) (17.5+15)+(17.5+20)=17.5(15+20)=17.5*35